Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2024 logbook, Contact: F.Wolff-Fabris, A.Galler, H.Sinn  Not logged in ELOG logo
Message ID: 103     Entry time: 13 May 2024, 08:48
Author: Ulrike Boesenberg 
Group: MID 
Subject: Shift summary 

Weekend summary:

Overall stable beam delivery with 300bunches @18keV and 750uJ

-issue with XGM on Saturday night/morning

-The intensity in SA2 over the train is dropping of when SA3 is changing some parameters. BKR can fix it within a couple of minutes

-The users found suitable working conditions to overcome the limitations from the 10bar He supply line.

ELOG V3.1.4-7c3fd00